DragonFlights, Inc. is planning on purchasing a new flying dragon for their new route to Volantis. The cost of the dragon isWhat is the cost of capital (WACC) for DragonFlights?What is depreciation per year using straight line depreciation over 10 years and zero salvage value?
Answer #1

Answer : Calculation of WACC

: (a.) Calculation of Firm's After tax cost of Debt :

Using Financial Calculator

=RATE(nper,pmt,pv,fv)

where nper is Number of years to maturity i.e 20 * 2 = 40 (As coupons are paid semiannually therefore multiplied by 2)

pmt is Interest payment i.e 1000 * 14% = 140 / 2 = 70 (Divided by 2 As coupons are paid semiannually)

pv is Current Market Price

= 1250

Note : pv should be taken as negative.

fv is face value i.e 1000

=RATE(40,70,-1250,1000)

therefore , Cost of Debt is 5.45184%(Semiannual)

Before Tax Cost of Debt is 5.45184% * 2 = 10.90367%(Annual)

After Tax Cost of Debt = 10.90367% * (1 - 0.25) = 8.18%

(b.) Cost of Equity = Risk Free rate + (Beta * Market Risk Premium)

= 2.5% + (2.8 * 7%)

= 22.1%

(c.) Calculation of WACC :

Market Value of Equity = Number of shares * Market Price per share

=1 million * 116

= 116 million

Market Value of Debt = Number of Bonds * Market Price

= 50000 * 1250

= 62,500,000 or 62.5 million

Total Market Value = 116 million + 62.5 million

= 178.5 million

Weight of Equity = 116 / 178.5

= 0.64985994397 or 0.6499

Weight of Debt = 62.5 / 178.5

= 0.35014005602 or 0.3501

WACC = (Cost of after tax Debt * Weight of Debt) + (Cost of Equity * Weight of Equity)

= [ 8.18% * 0.3501] +[22.1% * 0.6499]

= 2.8634% + 14.3619%

=17.23%

Depreciation per year using SLM = Cost / Number of year of useful Life

= 10.3 million / 10

= 1.03 million